# Graph Adapter of EEG Foundation Models for Parameter Efficient Fine Tuning

### v1

@@ Introduction
In the MDD downstream task, the GAT-based EEG-GraphAdapter (EGA-GAT) achieved 2.7% and 12.8% higher F1-score and AUROC than the baseline BENDR model, respectively. However, the performance of EGA-GraphSAGE fell below the baseline, particularly in the F1-score. On the other hand, in the TUAB downstream task, EGA-GraphSAGE achieved 16.1% higher performance than the baseline in the F1-score. The GraphSAGE-based model consistently outperformed the baseline across all metrics, demonstrating stable and high model performance.
@@
The number of trainable parameters for each model (BENDR and three EGA models with two hidden layers with size 64) is shown in Table 4. Compared to fine-tuning the BENDR model with many parameters in a downstream task, it is expected that fine-tuning only EGA will be able to execute the task in about 1/6 to 1/4 of the calculation time. The number of parameters in GraphSAGE is approximately twice that of GCN and GAT due to the PyG implementation, where the parameter matrices for linear transformations are implemented separately for self-nodes and neighbor nodes in PyTorch Geometric.
@@
Theoretically, training time should be reduced in proportion to the number of model parameters, but the speedup was a maximum of about 17%. This is due to overheads other than model parameter updates, such as pre-fetching of EEG datasets. By parallelizing the data loading process, this performance bottleneck should be resolved, but this is not the scope of this research.
Related Work
EEG and Transformer-based Models
@@
Similar to the BENDR(Kostas, Aroca-Ouellette, and Rudzicz 2021) model, numerous methods have been proposed to pre-train Transformer-based models on EEG signals using self-supervised learning to enhance few-shot learning performance in downstream tasks. For instance, BrainWave (Yuan et al. 2024) introduces a Transformer-based foundation model that can be pre-trained on both EEG signals and more invasive but highly accurate intracranial electroencephalograph (iEEG) signals. MAEEG (Chien et al. 2022) also performs pre-training by masking and reconstructing embeddings derived from stacked convolutional layers, similar to BENDR. However, MAEEG uses a reconstruction loss instead of a contrastive loss, achieving more stable and accurate results on longer EEG signals than BENDR. EEG2Rep (Mohammadi Foumani et al. 2024) applies multiple masking techniques, such as semantic subsequence preserving (SSP), which consecutively masks semantically related parts of the data. This allows the model to learn more abstract features during pre-training, resulting in a robust pre-trained model resistant to noise in raw EEG data.
GNN Models in Spatial Representation
@@
In addition to Transformer-based models, several GNN-based models have been proposed to account for the relationships between EEG sensors. MGGCN (Sun et al. 2024) constructs a Brain Functional Connectivity Network (BFCN) based on the functional connections between sensors derived from EEG signals and uses a GCN-based model to predict Major Depressive Disorder (MDD). Similarly, EEG-GCNN (Wagh and Varatharajah 2020) represents the relationships between EEG sensors as a fully-connected graph with weights based on sensor distances and the correlations of time-series data using a GCN-based model for neurological disorder prediction. EEG-GNN (Demir et al. 2021) takes a similar approach by representing the sensor relationships as a weighted graph, with weights based on sensor distances. This model, using GraphSAGE or GIN, predicts subjects’ motor patterns.
@@
These GNN-based models represent the relationships between EEG sensors as a graph structure, learning to capture functional connectivity between different brain regions. However, except for EEG-GNN, the time-series EEG signals corresponding to the graph vertices (sensors) are typically simplified, focusing on frequency-domain intensity. This approach does not account for the specific short- and long-term signal variations that Transformer-based models are designed to capture.
PEFT and Graph Adapter
@@
Many methods have been proposed for PEFT that are designed to reduce the number of parameters. One typical example is Low-Rank Adaptation (LoRA)(Hu et al. 2022), which reduces the computation costs for fine-tuning by adding low-rank matrices to the back of the pre-trained model. In addition, several methods have been proposed to combine an LLM with a GNN as an adapter. This makes integrating the LLM’s semantic representation with the graph’s spatial information possible.
@@
For example, G-Adapter (Gui, Ye, and Xiao 2024) enhances molecular structure prediction by inserting parameters into the feed-forward network (FFN) of a pre-trained Transformer model, allowing for incorporating spatial information, such as adjacency matrices, during fine-tuning. Recently, PEFT techniques have gained traction in applying large-scale pre-trained models like LLMs to various applications. Specifically, methods have been proposed that combine LLMs and GNNs for tasks involving Text-Attributed Graphs (TAGs), where text information is attached to nodes in the graph, such as in Knowledge Graphs.
@@
One such method, TAPE (He et al. 2024), encodes textual features such as article titles in citation networks using a pre-trained LLM and fine-tunes only a small-scale language model and a GNN. This approach maintains the LLM’s semantic richness while improving performance on graph-structured classification tasks. Similarly, GraphAdapter (Huang et al. 2024) fixes the parameters of a pre-trained LLM during downstream tasks, updating only the GNN and a fusion layer that integrates both models, thus improving efficiency on prediction tasks in TAGs, such as citation networks and social networks.
@@
Conclusion and Future Work
In this study, we proposed EEG-GraphAdapter (EGA), a method designed to efficiently solve healthcare-related downstream tasks with EEG signals by introducing a GNN model that learns the relationships between EEG sensors. With a GNN-based module as an adapter, we efficiently utilized pre-trained models, such as BENDR, that capture the time-series representation of EEG signals. This allowed for efficient learning of EEG representations tailored to each downstream task with minimal computational cost. Through experiments on four downstream tasks, EGA demonstrated superior accuracy with shorter training times compared to models that fine-tuned only the BENDR model or fine-tuned both the BENDR and GNN models. These results highlight the effectiveness of EGA in achieving high performance with reduced computational requirements.
As for future work, we aim to evaluate the broader applicability of the proposed EGA across various pre-trained models and downstream tasks. While this study employed the convolution-based BENDR model as the pre-trained model, we will explore other Transformer-based models such as BrainWave (Brant-2). Given that Transformer-based models typically require substantial computational resources for fine-tuning, the graph adapter may prove even more effective in such contexts. Additionally, we plan to evaluate and verify the performance of EGA in other healthcare-related downstream tasks, where the relationship between EEG sensors is considered to have a significant impact on the prediction results. For instance, in autism spectrum disorder (ASD), weak functional connectivity between the frontal and temporal lobes or between the cerebral hemispheres has been observed. In contrast, abnormalities in functional connectivity between the prefrontal cortex and the striatum have been observed in attention deficit hyperactivity disorder (ADHD). Since the amount of EEG data sets related to neurological disorders is often limited, the effect of introducing EGA could be particularly pronounced in these cases.
Acknowledgments
This work is partially supported by JSPS KAKENHI Grant JP21K17749 and 23K28098.

### v2

@@ 3.2 Performance of EGA
In the MDD downstream task, the GAT-based EEG-GraphAdapter (EGA-GAT) achieved a 12.8% higher F1-score and a 2.7% higher AUROC than the baseline BENDR model, respectively. Meanwhile, the GCN-based EEG-GraphAdapter (EGA-GCN) also outperformed the baseline, indicating that both GAT and GCN can effectively capture spatial information for MDD. However, the performance of EGA-GraphSAGE fell below the baseline, particularly in the F1-score.
On the other hand, in the TUAB downstream task, EGA-GraphSAGE achieved a 16.1% improvement over the baseline in the F1-score. The GraphSAGE-based model consistently outperformed the baseline across all metrics, demonstrating stable and high model performance. Overall, these results suggest that the most suitable GNN architecture can vary depending on the downstream task, underscoring the importance of selecting an appropriate graph-based module for each application.
@@ 3.3 Effectiveness of EGA
The number of trainable parameters for each model (BENDR and three EGA models with two hidden layers with size 64) is shown in Table 4. Compared to fine-tuning the BENDR model with many parameters, fine-tuning only EGA is expected to complete the task in roughly 1/6 to 1/4 of the time. The number of parameters in GraphSAGE is approximately twice that of GCN and GAT due to the PyG implementation, where the parameter matrices for linear transformations are implemented separately for self-nodes and neighbor nodes in PyTorch Geometric.
@@
Theoretically, training time should be reduced in proportion to the number of model parameters, but the speedup was a maximum of about 17%. This is due to overheads other than model parameter updates, such as pre-fetching of EEG datasets. By parallelizing the data loading process, this performance bottleneck should be mitigated, but this is beyond the scope of our research.
4 Related Work
4.1 EEG and Transformer-based Models
@@
Similar to the BENDRbendr model, numerous methods have been proposed to pre-train Transformer-based models on EEG signals using self-supervised learning to enhance few-shot learning performance in downstream tasks. For instance, BrainWave brainwave introduces a Transformer-based foundation model that can be pre-trained on both EEG signals and more invasive but highly accurate intracranial electroencephalograph (iEEG) signals. MAEEG maeeg also performs pre-training by masking and reconstructing embeddings derived from stacked convolutional layers, similar to BENDR. However, MAEEG uses a reconstruction loss instead of a contrastive loss, achieving more stable and accurate results on longer EEG signals than BENDR. EEG2Rep eeg2rep applies multiple masking techniques, such as semantic subsequence preserving (SSP), which consecutively masks semantically related parts of the data. This allows the model to learn more abstract features during pre-training, resulting in a robust pre-trained model resistant to noise in raw EEG data.
4.2 GNN Models in Spatial Representation
@@
In addition to Transformer-based models, several GNN-based models have been proposed to account for the relationships between EEG sensors. MGGCN mggcn constructs a Brain Functional Connectivity Network (BFCN) based on the functional connections between sensors derived from EEG signals and uses a GCN-based model to predict Major Depressive Disorder (MDD). Similarly, EEG-GCNN eeg-gcnn represents the relationships between EEG sensors as a fully-connected graph with weights based on sensor distances and the correlations of time-series data using a GCN-based model for neurological disorder prediction. EEG-GNN eeg-gnn takes a similar approach by representing the sensor relationships as a weighted graph, with weights based on sensor distances. This model, using GraphSAGE or GIN, predicts subjects’ motor patterns.
@@
These GNN-based models represent the relationships between EEG sensors as a graph structure, learning to capture functional connectivity between different brain regions. However, except for EEG-GNN, the time-series EEG signals corresponding to graph vertices (sensors) are typically simplified to focus on frequency-domain intensity. This approach does not account for the specific short- and long-term signal variations that Transformer-based models are designed to capture.
4.3 PEFT and Graph Adapter
@@
Many methods have been proposed for PEFT that are designed to reduce the number of parameters. One typical example is Low-Rank Adaptation (LoRA)lora , which reduces the computation costs for fine-tuning by adding low-rank matrices to the top of the pre-trained model. In addition, several methods have been proposed to combine an LLM with a GNN as an adapter. This makes integrating the LLM’s semantic representation with the graph’s spatial information possible.
@@
For example, G-Adapter g-adapter enhances molecular structure prediction by inserting parameters into the feed-forward network (FFN) of a pre-trained Transformer model, allowing for incorporating spatial information, such as adjacency matrices, during fine-tuning. Recently, PEFT techniques have gained traction in applying large-scale pre-trained models like LLMs to various applications. Specifically, methods have been proposed that combine LLMs and GNNs for tasks involving Text-Attributed Graphs (TAGs), where text information is attached to nodes in the graph, such as in Knowledge Graphs.
@@
One such method, TAPE tape , encodes textual features such as article titles in citation networks using a pre-trained LLM and fine-tunes only a small-scale language model and a GNN. This approach maintains the LLM’s semantic richness while improving performance on graph-structured classification tasks. Similarly, GraphAdapter graphadapter fixes the parameters of a pre-trained LLM during downstream tasks, updating only the GNN and a fusion layer that integrates both models, thus improving efficiency on prediction tasks in TAGs, such as citation networks and social networks.
@@
5 Conclusion and Future Work
In this study, we proposed EEG-GraphAdapter (EGA), a method designed to efficiently solve healthcare-related downstream tasks with EEG signals by introducing a GNN model that learns the relationships between EEG sensors. With a GNN-based module as an adapter, we efficiently utilized pre-trained models, such as BENDR, that capture the time-series representation of EEG signals. This allowed for efficient learning of EEG representations tailored to each downstream task with minimal computational cost. Through experiments on the two downstream tasks, EGA demonstrated both superior accuracy and shorter training times compared to the baseline that fully fine-tuned BENDR. These results highlight the effectiveness of EGA in achieving high performance with reduced computational requirements.
As for future work, we aim to evaluate the broader applicability of the proposed EGA across various pre-trained models and downstream tasks. While this study employed the convolution-based BENDR model as the pre-trained model, we will explore other Transformer-based models such as BrainWave (Brant-2)brainwave . Given that Transformer-based models typically require substantial computational resources for fine-tuning, the graph adapter may prove even more effective in such contexts. Additionally, we plan to evaluate and verify the performance of EGA in other healthcare-related downstream tasks, where the relationship between EEG sensors is considered to have a significant impact on the prediction results. For instance, in autism spectrum disorder (ASD), weak functional connectivity between the frontal and temporal lobes or between the cerebral hemispheres has been observed. In contrast, abnormalities in functional connectivity between the prefrontal cortex and the striatum have been observed in attention deficit hyperactivity disorder (ADHD). Since the amount of EEG data sets related to neurological disorders is often limited, the effect of introducing EGA could be particularly pronounced in these cases. Therefore, EGA offers a promising approach for resource-limited clinical EEG applications, paving the way for more accurate and scalable neurological disorder detection.
Acknowledgements.
This work is partially supported by JSPS KAKENHI Grant JP21K17749 and JP23K28098.In this research work, we used the “mdx: a platform for building data-empowered society”.